how do you solve 1-5 using the substitution method.
1. 4x+3y=11
x=3y-1
2. 2x+3y=5
x=4-2y
3. 3x-2y=11
y=5-2x
4. x+y=-5
4x-2y=10
5. y=2x-1
y=2-x

Answer by unlockmath(1688)

Hello,
I'll do the first one, let's see if you can do the rest of them.
4x+3y=11
x=3y-1
Substitute the second equation into the first like this:
4(3y-1)+3y=11
Expand:
12y-4+3y=11
Add 4 and combine:
15y=15
Divide by 15:
y=1
Plug this into either equation to find x:
x=2
Make sense?
RJ
www.math-unlock.com